# Synchronization in a small-world network of non-identical Chialvo neurons

### v1

@@ Abstract
We investigate the synchronization dynamics of a neuron network constructed using the small-world algorithm. The stochastic version of the map-based Chialvo neuron model is used to simulate each node of the network. To represent non-identical neurons, we introduce a mismatch in one of the main model parameters. Our study explores the impact of this mismatch, noise intensity in the stochastic model, and coupling strength between neurons on synchronization and firing frequency. In our initial approach, we consider electrically coupled neurons with both excitatory and inhibitory connections. We have identified critical values of noise intensity, parameter mismatch, and rewiring probability that lead to effective synchronization within the network. Additionally, we observe that the balance between excitatory and inhibitory connections significantly influences global synchronization. Our findings provide insights into the mechanisms underlying synchronization dynamics in complex neuron networks.
@@ I Introduction
The phenomenon of collective synchronization occurs when a system of coupled oscillators spontaneously locks to a common behavior independently of its individual characteristics. This phenomenon has been observed in many different scientific fields such as biology [1, 2, 3], engineering [4, 5] and ecology [6, 7]. In addition to all these examples, the synchronization of brain neurons is acquiring special interest in neuroscience because it is a fundamental neural mechanism [8, 9]. Different types of synchronous behaviors between neurons have been described, including complete synchronization, phase synchronization, lag synchronization [10] and have been related with some of the brain functions. Thus, the transition between in-phase and anti-phase synchronization is related with the locomotor brain rhythms [11, 12] and the phase synchronization with neural integration and working memory [13, 14, 15]. On the other hand, pathological cases of synchronization may be a consequence of serious brain disorders, such as essential tremor, Parkinson disease, and epilepsy [16]. Moreover the synchronization in the brain network, the firing frequency of the neurons of the brain is a characteristic with particular interest because is related with the neural coding. Recent studies have proved that the timing of spikes plays a critical role in encoding, representing, and processing knowledge and events in the brain. [17, 18].
With all these considerations in mind, various models have been proposed to grasp the synchronization of coupled neurons. Among them are continuous neuron models based on ordinary differential equations, such as Hodgkin-Huxley [19, 20], FitzHugh-Nagumo [21], Morris-Lecar [22], and Hindmarsh-Rose [23, 24], which provide a realistic depiction of the electrophysiological processes within neurons. Conversely, discrete neuron models like Rulkov, Izhikevich, Courbage-Nekorkin-Vdovin [25, 26, 27, 28, 29] offer computational ease due to their simplicity, thereby reducing the computational workload.
In our work, we propose using the stochastic version of the Chialvo map [30] to understand synchronization in a neuron network. The effect of noise in excitable systems has been extensively studied for white and coloured noise [31, 32] and it has been found that the noise intensity is related with changes in the particular behavior of the neuron models [33, 34, 35]. Moreover, it is also known that some factors such as coupling strength [36, 37], coupling forms [38, 39], and diversity [40], play a key role in the global synchronization of the network. For all these reasons, in this work we will represent the diversity in the network introducing a parameter mismatch in the neuron model and will consider different types of coupling either excitatory or inhibitory. The global synchronization of the network and the firing frequency will be studied using numerical methods for different characteristics of the neurons and of the network itself. We have found critical values of the noise intensity that enhance the synchronization of the neuron network. We have observed that the presence of inhibitory couplings desynchronizes the system while the increase of the rewiring probability improves the stability of the neuron network. All these features are explained along the manuscript.
The organization of this paper is as follows. In Section II, we describe the stochastic Chialvo model. In Section III, we analyze the synchronization and the firing frequency in the small-world neuron network as a function of the mismatch between the neurons in one of the parameters of the model, the noise intensity an the proportion between the excitatory and the inhibitory couplings. The main conclusions of our results are provided in Section IV.
@@ II The stochastic Chialvo neuron model
To construct the neuron network, we will use the stochastic version of the map-based Chialvo model [30]. In this model, the variable $x$ is related to an instantaneous membrane potential of the neuron, while the variable $y$ stands for the recovery current. The parameter $I$ models the action of the ion current injected into the neuron, the parameter $a$ is related to the time of recovery ($a<1$), the activation-dependence of the recovery process is defined by the parameter $b$ ($b<1$) with the offset value $c$. In our case, the stochastic model, random disturbance $\varepsilon\xi_{t}$ is added in the parameter $I$ of the acting ion current. The parameter $\varepsilon$ represents the intensity of the noise, while $\xi_{t}$ is uncorrelated white Gaussian noise with parameters $\langle\xi_{t}\rangle=0,\;\langle\xi^{2}_{t}\rangle=1$. All this information is summarized in Eq. 1 which reads:
@@
This neuron model is capable of exhibiting rich dynamics and can adequately reproduce some of the basic features behind the firing dynamics of the neurons [30, 41]. For example, by modifying the parameter $b$, one could obtain different dynamical behaviors: fixed point, periodic, quasi-periodic or chaotic dynamics. These basic dynamics of the neuron can be appreciated in Fig. 1(a-d) where the potential action, $x_{t}$, is represented for some typical values of the model parameters ($a$, $c$ and $I$) and for different values of the parameter $b$.
@@ III Small-world neuron network
We propose to study the effect of the noise intensity and the mismatch in the parameter $b$ on the synchronization and the firing frequency in a neuron network. We will consider two different kinds of couplings between the neurons of the network: excitatory and inhibitory. With this purpose in mind, the first step is to construct the neuron network, that is, to connect mathematically a group of neurons. In our case, we connect the neurons in the well-known small-world network [42]. To build a small-world network of $N$ neurons, the first step is to order the $N$ neurons, and then connect each neuron with its $l$ first neighbors in a cyclical way. In this sense, the first node should be connected with the $l$ latest nodes. Finally, each node has a probability $p$ to rewire one of these edges with any other node. Thus, this kind of graph is characterized by three parameters: (i) $N$, the number of nodes; (ii) $l$, the number of first neighbors connected with each neuron; and (iii) $p$, the probability of rewiring an edge. This last parameter indicates the random nature of the graph, i.e., as $p$ increases the graph will be more random [42]. After the random rewiring of the network connections, each neuron will have a total number of connections $N_{i}$ that can change from one neuron to another. This parameter, $N_{i}$, plays an important role in the network, because the connection strength between the neuron $i$ and the neuron $j$ will be done by ($k/N_{i}$).
@@
Once we have developed the mathematical model for the neuron network (Eq. 3), we proceed to the numerical simulations. In our case, we will set the mean value of the parameter $b$ at $b=0.35$ and will introduce a random mismatch in the parameter $b$ of some neurons. The maximum value of this mismatch would be (±$\pm 0.01b$). We introduce some small differences between the neurons using this value, while still considering them similar.
It would be senseless to consider neurons located in the same brain region and part of the same local network as very different. Introducing a mismatch in the parameter $b$ for every neuron and the fact that some of the links between the neurons are inhibitory connections introduce some instability in the neuron network, significantly affecting the synchronization of the network. Previous simulations have led us to reject values of $b$ corresponding to the Region I of the Fig. 2, because in this case, the neuron is highly sensitive to the value of $b$ and even the small mismatch we propose makes almost impossible to have a great synchronization in the network without considering a very high coupling strength between the neurons, which would be unrealistic. We also discard values from the Region II because the results are quite similar to those obtained in the case we propose ($b=0.35$).
@@
The objective of the numerical simulations is to see how variations in the rest of the model parameters affect the global synchronization of the network and to the firing frequency of the neurons. To analize the synchronization between the neurons of the network, we introduce the order parameter $R$ as in Ref. [43]. This parameter allows us to measure the synchronization between different signals. In our case, we will apply this parameter to study the synchronization between the variable $x$ of each neuron from the network. Its mathematical expression is given by:
@@ III.0.1 Excitatory coupling in the small-world network
Figure 9: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where all neurons have the same values for the parameters, $b=0.35$, $a=0.89$, $c=0.28$ and $I=0.03$.
Figure 10: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where 25 out of the 50 neurons have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$.
Figure 11: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where all the 50 neurons of the network have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$.
@@ III.0.2 Inhibitory coupling in the small-world network
Normally, when two neurons are connected by inhibitory coupling, their synchronization worsens. But can this observation be extended to a neural network? Can we assume that the introduction of some inhibitory connections must desynchronize the system? To answer these questions, we propose to introduce a percentage of inhibitory connections in the network and analyze its effect on global synchronization. The percentage of inhibitory coupling has a great dependence on the area of the brain analyzed and on the type of neurons connected [44], and its effect is quite important in the global behavior of the neuron network [45, 46]. As in the previous section, the figures obtained for the cases corresponding to values of $b$ corresponding to the Regions II and III remarked in Fig. 1 are quite similar, so we focus our simulations on the case $b=0.35$. The procedure is similar to the excitatory case: we will start by considering that all the neurons have the same value for the parameter $b$ and gradually increase the number of neurons with a mismatch in this parameter.
@@
Figure 15: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where all neurons have the same values for the parameters, $b=0.35$, $a=0.89$, $c=0.28$ and $I=0.03$. In the network an $1\%$ of the connections are inhibitory couplings.
Figure 16: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where 25 out of the 50 neurons have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network an $1\%$ of the connections are inhibitory couplings.
Figure 17: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where all the 50 neurons from the network have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network an $1\%$ of the connections are inhibitory couplings.
@@
Figure 21: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where all neurons have the same values for the parameters, $b=0.35$, $a=0.89$, $c=0.28$ and $I=0.03$. In the network a $5\%$ of the connections are inhibitory couplings.
Figure 22: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where 25 out of the 50 neurons have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network a $5\%$ of the connections are inhibitory couplings.
Figure 23: Plot of the synchronization order parameter R. Representation of the mean value of the $ISI$ for a network where all the 50 neurons from the network have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network a $5\%$ of the connections are inhibitory couplings.
@@ IV Conclusions
We have investigated the global behavior of a neuron network organized in the well-known small-world pattern. We have used a stochastic version of the map-based Chialvo neuron model, where a random mismatch in the parameter $b$ was introduced in some neurons to induce inhomogeneity within the network. Under these conditions, we have analyzed the synchronization of the network and the inter-spike interval ($ISI$) as a function of the noise intensity $\varepsilon$ introduced in the stochastic model, the coupling strength between neurons $k$, and the mismatch in the parameter $b$.
Additionally, we examined the network’s global behavior based on its structure, by increasing the rewiring probability of network connections $p$ and adjusting the proportion between excitatory and inhibitory connections.
In general, synchronization is more pronounced in homogeneous networks compared to inhomogeneous ones; as the number of neurons with a mismatch increases, the system becomes more desynchronized. When the network consists solely of excitatory connections, synchronization is high and increases with coupling strength. However, transitioning some connections to inhibitory decreases the region of high synchronization in the parameter plane ($k$,$\varepsilon$) as the proportion of inhibitory connections rises. We have observed that increasing the rewiring probability $p$ enhances network synchronization, indicating that randomness in network connections adds stability to the system. Moreover, certain values of noise intensity optimize network synchronization, remaining almost constant regardless of coupling strength and the proportion of excitatory/inhibitory connections.
@@
Acknowledgments JU, JMS and MAFS acknowledge financial support from the Spanish State Research Agency (AEI) and the European Regional Development Fund (ERDF, EU) under Project No. PID2019-105554GB-I00 (MCIN/AEI/10.13039/501100011033). IB and LR acknowledge financial support from the Ministry of Science and Higher Education of the Russian Federation (project ”Ural Mathematical Center” N 075-02-2024-1428).

### v3

# Impact of Network Heterogeneity on Neuronal Synchronization

@@ Abstract
Synchronization dynamics is a phenomenon of great interest in many fields of science. One of the most important fields is neuron dynamics, as synchronization in certain regions of the brain is related to some of the most common mental illnesses. To study the impact of the network heterogeneity in the neuronal synchronization, we analyze a small-world network of non-identical Chialvo neurons that are electrically coupled. We introduce a mismatch in one of the model parameters to introduce the heterogeneity of the network. Our study examines the effects of this parameter mismatch, the noise intensity in the stochastic model, and the coupling strength between neurons on synchronization and firing frequency. We have identified critical values of noise intensity, parameter mismatch, and rewiring probability that facilitate effective synchronization within the network. Furthermore, we observe that the balance between excitatory and inhibitory connections plays a crucial role in achieving global synchronization. Our findings offer insights into the mechanisms driving synchronization dynamics in complex neuron networks.
@@ I Introduction
The phenomenon of collective synchronization occurs when a system of coupled oscillators spontaneously locks to a common behavior, independently of their individual characteristics. This phenomenon has been observed in many different scientific fields such as biology Peskin75 ; Liu97 ; aldridge99 , engineering Kourtchatov95 ; York91 , and ecology Buck76 ; Walker69 . In addition to all these examples, the synchronization of brain neurons is acquiring special interest in neuroscience because it is a fundamental neural mechanism Engel01 ; Uhlhaas06 . Different types of synchronous behaviors between neurons have been described, including complete synchronization, phase synchronization, and lag synchronization Boccaletti2018 . These behaviors have been related to various brain functions. For instance, the transition between in-phase and anti-phase synchronization is associated with locomotor brain rhythms Jia18 ; Song12 , while phase synchronization is related to neural integration and working memory Bambi00 ; Batista07 ; Varela01 . On the other hand, pathological cases of synchronization may result in serious brain disorders, such as essential tremor, Parkinson’s disease, and epilepsy Pare90 . Beyond synchronization in the brain network, the firing frequency of neurons in the brain is of particular interest because it relates to neural coding. Recent studies have shown that the timing of spikes plays a critical role in encoding, representing, and processing knowledge and events in the brain Fujii96 ; MacLeod96 . With all these considerations in mind, various models have been proposed to understand the synchronization of coupled neurons. Among them are continuous neuron models based on ordinary differential equations, such as Hodgkin-Huxley Hodgkin1952 ; Hodgkin1952_2 , FitzHugh-Nagumo FitzHugh1961 , Morris-Lecar morris:lecar , and Hindmarsh-Rose Hindmarsh1982 ; Shilnikov2008 , which provide a realistic depiction of the electrophysiological processes within neurons. Conversely, discrete neuron models like Rulkov, Izhikevich, and Courbage-Nekorkin-Vdovin Rulkov01 ; Ibarz2011 ; Gir2013 ; KT96 ; Courbage2007 offer computational ease due to their simplicity, thereby reducing the computational workload. In our work, we propose using the stochastic version of the Chialvo map Chialvo95 to understand synchronization in a neuron network. The effect of noise in excitable systems has been extensively studied for both white and colored noise zambrano:2010 ; LinGar04 , and it has been found that noise intensity is related to changes in the particular behavior of neuron models bruss:2023 ; BNR18 ; BNR_CSF18 . Moreover, it is also known that factors such as coupling strength Sun15 ; Shen08 , coupling forms Wang2011 ; Sriram2023 , and diversity Tang11 play key roles in the global synchronization of the network. The choice of the stochastic Chialvo model over other models is motivated by the observation that critical values of noise and coupling significantly enhance network synchronization—a phenomenon not observed in other models in the same way as in this study. Some of those features have been studied in recent works. In Ref. anjana24 the authors studied the dynamics of the system in a simple network configuration with identical neurons, while in Ref. Kuznetsov24 the authors described the dynamics of the system when two or three non-identical neurons are connected. In this work, we take into account the heterogeneity of the network by introducing a parameter mismatch in the neuron model and consider different types of coupling, either excitatory or inhibitory but in a much more complicated network structure. The global synchronization of the network and the firing frequency will be studied using numerical methods for various characteristics of the neurons and the network itself. We have found critical values of noise intensity that enhance the synchronization of the neuron network. We observed that the presence of inhibitory couplings desynchronizes the system, while an increase in the rewiring probability improves the stability of the neuron network. All these features are explained in detail throughout the manuscript. The organization of this paper is as follows. In Section II, we describe the stochastic Chialvo model. In Section III, we analyze the synchronization and the firing frequency in the small-world neuron network as a function of the mismatch between the neurons in one of the parameters of the model, the noise intensity an the proportion between the excitatory and the inhibitory couplings. The main conclusions of our results are provided in Section IV.
@@ II The stochastic Chialvo neuron model
To construct the neuron network, we will use the stochastic version of the map-based Chialvo model Used24 . This model is based in the deterministic Chialvo model Chialvo95 where the variable $x$ is related to an instantaneous membrane potential of the neuron, while the variable $y$ stands for the recovery current. In this case a random disturbance $\varepsilon\xi_{t}$ is added in the parameter $I$ of the acting ion current injected into the neuron. The parameter $\varepsilon$ represents the intensity of the noise, while $\xi_{t}$ is uncorrelated white Gaussian noise with parameters $\langle\xi_{t}\rangle=0,\;\langle\xi^{2}_{t}\rangle=1$. The parameter $a$ is related to the time of recovery ($a<1$), the activation-dependence of the recovery process is defined by the parameter $b$ ($b<1$) with the offset value $c$. All this information is summarized in Eq. 1:
@@
This neuron model is capable of reproducing some of the basic features behind the firing dynamics of the neurons Chialvo95 ; Hoppensteadt86 . For example, by modifying the parameter $b$, one could obtain different dynamical behaviors: fixed point, periodic, quasi-periodic or chaotic dynamics. These basic dynamics of the neuron can be appreciated in Fig. 1(a-d) where the potential action, $x_{t}$, is represented for some typical values of the model parameters ($a$, $c$ and $I$) and for different values of the parameter $b$.
@@ III Small-world neuron network
In a previous work Used24 , the effect of noise intensity and the mismatch in the parameter $b$ on synchronization and the firing frequency was anlyzed in a system constructed by connecting only two non-identical neurons. In this simple model we also considered two different types of couplings between the neurons: excitatory and inhibitory. To generalize the results obtained in this previous work we propose a similar study but in a small-world neuron network. With this purpose in mind, the first step is to construct the neuron network, that is, to connect mathematically a group of neurons. In our case, we connect the neurons in the well-known small-world network Watts98 . To build a small-world network of $N$ neurons, the first step is to order the $N$ neurons, and then connect each neuron with its $l$ first neighbors in a cyclical way. In this sense, the first node should be connected with the $l$ latest nodes. Finally, each node has a probability $p$ to rewire one of these edges with any other node. Thus, this kind of graph is characterized by three parameters: (i) $N$, the number of nodes; (ii) $l$, the number of first neighbors connected with each neuron; and (iii) $p$, the probability of rewiring an edge. This last parameter indicates the random nature of the graph, i.e., as $p$ increases the graph will be more random Watts98 . After the random rewiring of the network connections, each neuron will have a total number of connections $N_{i}$ that can change from one neuron to another. This parameter, $N_{i}$, plays an important role in the network, because the connection strength between the neuron $i$ and the neuron $j$ will be done by ($k/N_{i}$).
@@
Once we have developed the mathematical model for the neuron network (Eq. 3), we proceed to the numerical simulations. In our case, we will set the mean value of the parameter $b$ at $b=0.35$ and will introduce a random mismatch in the parameter $b$ of some neurons. The maximum value of this mismatch would be (±$\pm 0.01b$). We introduce some small differences between the neurons using this value, while still considering them similar. It would be senseless to consider neurons located in the same brain region and part of the same local network as very different. Introducing a mismatch in the parameter $b$ for every neuron and the fact that some of the links between the neurons are inhibitory connections introduce some instability in the neuron network, significantly affecting the synchronization of the network. Previous simulations have led us to reject values of $b$ corresponding to the Region I of the Fig. 2, because in this case, the neuron is highly sensitive to the value of $b$ and even the small mismatch we propose makes almost impossible to have a great synchronization in the network without considering a very high coupling strength between the neurons, which would be unrealistic. We also discard values from the Region II because the results are quite similar to those obtained in the case we propose ($b=0.35$).
@@
The objective of the numerical simulations is to see how variations in the rest of the model parameters affect the global synchronization of the network and to the firing frequency of the neurons. To analize the synchronization between the neurons of the network, we introduce the order parameter $R$ as in Ref. Ojalvo93 . This parameter allows us to measure the synchronization in phase and in amplitude between different signals. In our case, we will apply this parameter to study the synchronization between the variable $x$ of each neuron from the network. Its mathematical expression is given by:
@@ III.0.1 Excitatory coupling in the small-world network
Figure 9: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where all neurons have the same values for the parameters, $b=0.35$, $a=0.89$, $c=0.28$ and $I=0.03$.
Figure 10: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where 25 out of the 50 neurons have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$.
Figure 11: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where all the 50 neurons of the network have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$.
@@ III.0.2 Inhibitory coupling in the small-world network
Normally, when two neurons are connected by inhibitory coupling, their synchronization worsens. But can this observation be extended to a neural network? Can we assume that the introduction of some inhibitory connections must desynchronize the system? To answer these questions, we propose to introduce a percentage of inhibitory connections in the network and analyze its effect on global synchronization. The percentage of inhibitory coupling has a great dependence on the area of the brain analyzed and on the type of neurons connected attila99 , and its effect is quite important in the global behavior of the neuron network Borgers03 ; Borgers05 . As in the previous section, the figures obtained for the cases corresponding to values of $b$ corresponding to the Regions II and III remarked in Fig. 1 are quite similar, so we focus our simulations on the case $b=0.35$. The procedure is similar to the excitatory case: we will start by considering that all the neurons have the same value for the parameter $b$ and gradually increase the number of neurons with a mismatch in this parameter.
@@
Figure 15: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where all neurons have the same values for the parameters, $b=0.35$, $a=0.89$, $c=0.28$ and $I=0.03$. In the network an $1\%$ of the connections are inhibitory couplings.
Figure 16: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where 25 out of the 50 neurons have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network an $1\%$ of the connections are inhibitory couplings.
Figure 17: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where all the 50 neurons from the network have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network an $1\%$ of the connections are inhibitory couplings.
@@
Figure 21: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where all neurons have the same values for the parameters, $b=0.35$, $a=0.89$, $c=0.28$ and $I=0.03$. In the network a $5\%$ of the connections are inhibitory couplings.
Figure 22: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where 25 out of the 50 neurons have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network a $5\%$ of the connections are inhibitory couplings.
Figure 23: Plot of the inter-spike interval ($ISI$). Representation of the mean value of the $ISI$ for a network where all the 50 neurons from the network have a random value for the parameter $b$. The mean value of the parameter $b=0.35$ and the mismatch is randomly selected with maximum value $\Delta b=0.01b$. The rest of the parameters are equal for all the neurons, $a=0.89$, $c=0.28$ and $I=0.03$. In the network a $5\%$ of the connections are inhibitory couplings.
@@ IV Conclusions
We have numerically investigated the global behavior of a neuron network organized in the well-known small-world pattern. We used a stochastic version of the map-based Chialvo neuron model, introducing a random mismatch in the parameter $b$ in some neurons to induce heterogeneity within the network. Under these conditions, we numerically analyzed the synchronization of the network and the inter-spike interval ($ISI$) as a function of the noise intensity $\varepsilon$ introduced in the stochastic model, the coupling strength between neurons $k$, and the mismatch in the parameter $b$. Additionally, we examined the network’s global behavior based on its structure by increasing the rewiring probability of network connections $p$ and adjusting the proportion between excitatory and inhibitory connections. In general, synchronization is more pronounced in homogeneous networks compared to heterogeneous ones. As the number of neurons with parameter mismatches increases, the system becomes more desynchronized. When the network consists solely of excitatory connections, synchronization is high and further increases with coupling strength. However, transitioning some connections to inhibitory decreases the region of high synchronization in the parameter plane ($k$,$\varepsilon$) as the proportion of inhibitory connections rises.
The most significant findings relate to both the probability of rewiring and the intensity of noise, with the latter being particularly striking, as it suggests the existence of noise levels that substantially enhance system synchronization. Regarding rewiring probability, we observe that increasing $p$ improves network synchronization. This effect remains consistent across all conditions, regardless of the proportion of neurons with mismatches or the specific types of connections within the network. Thus, we deduce that introducing randomness into the network enhances its stability. Regarding the impact of noise on network synchronization, we identify specific noise values that lead to a marked increase in synchronization. These values are approximately centered at $\varepsilon=0.8\cdot 10^{-3}$ and $\varepsilon=2\cdot 10^{-3}$ and remain nearly constant regardless of coupling strength or the proportion of excitatory/inhibitory connections. When all connections are excitatory, the improvement in network synchronization at these noise levels is particularly evident, especially at lower rewiring probabilities. As the rewiring probability increases, the effect of noise diminishes, as the system reaches high synchronization levels. The influence of these noise levels becomes more pronounced when inhibitory connections are introduced. In both cases, optimal synchronization occurs at the aforementioned noise levels. This effect is noticeable with $1\%$ inhibitory connections but significantly more pronounced with $5\%$ inhibitory connections.
@@
Acknowledgments JU, JMS and MAFS acknowledge financial support from the Spanish State Research Agency (AEI) and the European Regional Development Fund (ERDF, EU) under Project Nos. PID2019-105554GB-I00 and  PID2023-148160NB-I00 (MCIN/AEI/10.13039/501100011033). IB and LR acknowledge financial support from the Ministry of Science and Higher Education of the Russian Federation (project “Ural Mathematical Center”, Agreement N 075-02-2025-1719/1).